# Statistical properties of Markov shifts (part I)

### v1

@@ 3.1.2. A truncation argument
Let Assumption 2.7 hold with $a<\infty$ and let $b$ be defined by $1/3=1/a+1/b$ (note that $b=\frac{3a}{a-3}$). Let $\zeta,\varepsilon,c_{0},c_{1}$ be like in that assumption. Let us fix some $M>0$. Define a function $G_{M}:{\mathbb{R}}\to{\mathbb{R}}$ as follows. Set $G_{M}(x)=x$ if $|x|\leq M$, set $G_{m}(x)=0$ if $|x|\geq 2M$ and on [2M,M][-2M,-M],  let $G_{M}$ identify with the linear function connecting (2M,0)(-2M,0) and (M,M)(-M,-M), while on $[M,2M]$ let it identify with the linear function connecting $(2M,0)$ and $(M,M)$. Then
@@ 3.2.2. A truncation argument
where $1/3=1/p+1/a$ for $a\leq p$. Therefore, with $\tilde{S}_{n}=\sum_{j=1}^{k_{n}}\tilde{Y}_{j,n}$ we have
(3.10) $\|S_{n}-\tilde{S}_{n}\|_{L^{3}}\leq C(1+\sigma_{n}^{2}M_{n}^{-p/a})=O(1)$
@@ 3.4. Reduction of optimal CLT rates from two sided functionals to one sided ones
Finally to deduce Theorem 2.22 for $S_{n}f$ from the corresponding result for $S_{n}g$ let us take some b<a1b<a-1. By Theorem 2.22, we can couple $S_{n}g$ with a standard normal random variable $Z$ so that $\|S_{n}g/{\sigma}_{n}-Z\|_{L^{b}}\leq C{\sigma}_{n}^{-(1-u)}$. Now by Berkes–Philipp Lemma [5, Lemma A.1], we can also couple all three random variables $S_{n}g$, $S_{n}f$ and $Z$ so that (3.17) still holds under the new probability law. ∎
@@ 4.1. Products of positive matrices and other operators
We assume that there are (possibly random) Birkhoff cones ${\mathcal{C}}_{j}\subset B_{j}$ and $n_{0}\in{\mathbb{N}}$ such that for all $j$ we have $A(X_{j}){\mathcal{C}}_{j}\subset{\mathcal{C}}_{j+1}$ (almost surely) and the projective diameter of $\textbf{A}_{j}^{n_{0}}{\mathcal{C}}_{j}$ inside ${\mathcal{C}}_{j+n_{0}}$ does not exceed some constant $d_{0}<\infty$ which is independent of $j$ (see [38, Appendix A] for an overview of projective metrics and cones). Finally, let us assume that the cones are regenerating in the sense of [62, Section 5], namely there exist $r\in{\mathbb{N}}$ and $C>0$ such that every $g\in B_{j}$ can be written as
@@
Each $A_{j}(X_{j})$ is the random transfer operator associated with a random expanding map $T_{X_{j}}$ satisfying the conditions of [38, Ch.5]. That is
@@
for a Hölder continuous random functions $\phi_{X_{j}}$ (which are uniformly Hölder continuous) and Hölder continuous functions $g$. Then by [38, Theorem] there are random cones satisfying the above conditions and (4.1) holds.
@@ 4.4.1. Processes with initial condition
First, by [19, Corollary 5.3] we have
@@
where the second inequality uses the Hölder inequality. Notice next that by applying again [19, Corollary 5.3] we see that
@@
for all $s\geq 0$ and measurable sets $A\in\sigma\{X_{j}:j\leq s\}$ and $B\in{\sigma}\{X_{j}:j>s\}$. Note that in the notations of [9, Eq. (1.6) and Eq. (2.2)] we have $\psi_{U}(1)=\psi^{*}(1)-1$.
@@ 4.4.2. Processes without initial condition
Now, by applying again [19, Corollary 5.3] we see that
@@ 4.5. Application to $\text{GARCH}(\mathfrak{p},\mathfrak{q})$ sequences
with $\mu,\alpha_{i},\beta_{j}\in{\mathbb{R}},\mu>0$. We refer to [35, Example 3.5] for more references and motivation for considering such processes. We assume here that with $r=\max(\mathfrak{p},\mathfrak{q})$,
@@
Now, as explained in [35, Example 3.5] we have
@@
Arguing like in [35, Example 3.5] one can show that
@@ 4.8. Subshifts of finite type and Hölder on average observables
Another example which is relevant to our setup is when working with measure of maximal entropy of a subshift of finite type $T$. Let us briefly recall the definition of a subshift of finite type. Let $\mathcal{A}$} be a finite set and let $(A_{i,j})_{i,j\in{\mathcal{A}}}$ be a matrix with 010-1 entries such that $A^{M}$ has only positive entries for some $M$. Let $\Sigma=\{(x_{i})\in{\mathcal{A}}^{{\mathbb{N}}}:A_{x_{i},x_{i+1}}=1\}$ and let $\sigma:\Sigma\to\Sigma$ be the left shift. Let $\mu$ be the unique measure of maximal entropy (see [6]). Then (see [6]) when viewed as random variables $(X_{j})_{j\geq 0}$ whose path is distributed according to $\mu$ the coordinates $X_{j}$ form a $\psi$-mixing Markov chain, and so we can prove limit theorems for partial sums of the form $\sum_{j=0}^{n-1}f_{j}\circ T^{j}$ for functions $f_{j}$ which are only Hölder continuous on average, that is under the assumption that
@@
for some $s$ and $\delta$. Indeed, Assumption 2.6 is in force. In fact, even when $f_{j}=f$ does not depend on $j$ these results seem to be new. Moreover, we can also consider non-stationary SFT $(T_{j})$ like in [24] since also in that case the coordinates are $\psi$-mixing (see [39]).
@@
When considering non stationary SFT (see [24]) we can get results for Markov measures, and so also for small perturbations of Anosov maps, see [24, Appendix C].
@@ 5.4. A complex Perron Frobenius theorem
Under one of Assumptions 2.6, 2.7 and 2.10 denote $B_{j}={\mathcal{B}}_{j,q,p,\delta^{\prime}}$ (where under Assumption 2.6 we have $\delta^{\prime}=\delta$). Under Assumption 2.11 denote $B_{j}={\mathcal{B}}_{j,\infty,\infty,\delta^{1/2}}$. Next, by applying [24, Theorem D.2] we get the following corollary of Theorem 5.1.
@@ 6.1. A martingale coboundary representations, the asymptotic behavior of the variance
Finally, the proof of the last statement proceeds like the proof of [3, Theorem 3.5], with minor modifications. ∎
@@ 6.2. Quadratic variation and moment estimates
First, to simplify the notation let us assume that $j=0$. The argument below is similar to the first part of the proof of [15, Theorem 4.1]. First, we write
@@
for some constant $C_{q}>0$. Recall the following version of Burkholder’s inequality for martingales (see [54, Theorem 2.12]). Let $\mathfrak{d}_{1},....,\mathfrak{d}_{n}$ be a martingale difference with respect to a filtration $({\mathcal{G}}_{j})_{j=1}^{n}$ on a probability space. Let $D_{n}=\mathfrak{d}_{1}+\mathfrak{d}_{2}+...+\mathfrak{d}_{n}$ and $E_{n}=\mathfrak{d}_{1}^{2}+\mathfrak{d}_{2}^{2}+...+\mathfrak{d}_{n}^{2}$. Then, for every $s\geq 2$ there are constants $c_{s},C_{s}>0$ depending only on $s$ such that
@@ 6.6. The sequential pressure function and its approximation properties
There exist $r_{0}>0,C_{0}>0$ and $n_{0}$ such that on $t\in[-r_{0},r_{0}]$ for all $n\geq n_{0}$ we can develop a branch $\Lambda_{j,n}(t)$ of $\ln\mu_{j}(e^{itS_{j,n}g})={\mathbb{E}}[e^{itS_{j,n}g}]$ such that for $s=0,1,2,3$ we have
@@
Notice that the first $k$ derivatives of $t\to\bar{E}_{j}^{t,n}$ are also of order $O(\delta_{2}^{n})$ (by the differentiation rule for derivatives of products). Hence for $s=0,1,2,3$ we have
@@
for some constant $C^{\prime}_{k}$. Since $k\geq 3$ we see that [40, Assumption ] is met. Thus Theorems 2.21 and 2.22 follows by [40, Theorem ] and [40, Theorem].
@@
As before, by Proposition 3.9 it is enough to prove both theorems for $S_{n}g$ (or $\tilde{S}_{n}$ in the notations of Section 3.2). Let $\Lambda_{j,n}(t)$ be the branch of $\ln{\mathbb{E}}[e^{itS_{j,n}g}]$ from Lemma 6.5. Note also that by rescaling under one of Assumptions 2.7 or 2.10 we can always assume that $u=0$ in Theorems 2.21 and 2.22. Recall that we assumed that ${\mathbb{E}}[f_{j}]=0$ which implies that ${\mathbb{E}}[g_{j}]=0$. By applying again [40, Theorem 5] and [40, Theorem 9] and [40, Corollary 11]. Theorems 2.21 and 2.22 for $S_{n}g$ will follow from the following result.
@@
In the circumstances of Theorems 2.21 and 2.22 there are constants ${\delta}_{3}>0$ and $C_{3}>0$ such that for all $s=0,1,2,3$ we have
(6.11) $\sup_{t\in[-{\delta}_{3},{\delta}_{3}]}|\Lambda^{(s)}_{0,n}(t)|\leq C_{3}{\sigma}_{n}^{2}.$
@@
Notice that it is enough to prove (6.11) with $s=3$. Indeed, if $|t|\leq\delta_{3}$ then
$\Lambda^{{}^{\prime\prime}}_{0,n}(t)=\int_{0}^{t}\Lambda^{{}^{\prime\prime\prime}}_{0,n}(x)dx=O({\sigma}_{n}^{2})$
and similarly $\Lambda^{(s)}_{0,n}(t)=O({\sigma}_{n}^{2})$ for $s=0,1$.
@@
Applying [23, Lemma 43] with $S=S_{j,n}g=S_{j,n}g-{\mathbb{E}}[S_{j,n}g]$ we see that there is $r=r(B)$ such that if $t\in[-r,r]$ then
(6.13) $|\Lambda_{j,n}^{(3)}(t)|\leq D_{0}{\mathbb{E}}[|S|^{3}]$
for some constant $D_{0}$. Now the result follows since $\|S_{j,n}f-S_{j,n}g\|_{L^{3}}\leq C$ for some constant $C$ (see (3.10), (3.6) and Lemma 3.5). ∎
@@
$\sup_{t\in[-r_{0},r_{0}]}\left|\Pi_{I_{j}}^{(3)}(t)\right|\leq D_{0}(1+{\mathbb{E}}[|S_{I_{j}}f|^{3}]).$
@@
(6.16) $\sup_{t\in[-r_{k},r_{k}]}\left|\Pi_{0,n}^{(3)}(t)\right|\leq D_{0}k_{n}+\sum_{j=1}^{k_{n}}{\mathbb{E}}[|S_{I_{j}}f|^{3}]=O({\sigma}_{n}^{2})$
@@
$\sup_{t\in[-r_{0},r_{0}]}\left|\Lambda_{n}^{(3)}(t)\right|\leq\tilde{D}{\sigma}_{n}^{2}$
@@ 7. Large deviations
where $g_{j}$ are the functions from Lemma 3.5. Denote $B_{j}={\mathcal{B}}_{j,\infty,\infty,\delta^{1/2}}$. Then, since $\sup_{j}\|g_{j}\|_{j,\infty,\infty,\delta^{1/2}}<\infty$ we see that $L_{j,z}$ are uniformly analytic in $z$. By applying [24, Theorem D.2] we get the following corollary of Theorem 5.1.
@@
Now, notice that (see the arguments in [38, Ch.5]) $\int\ln{\lambda}_{\omega}(t)d\mathbb{P}(\omega)=1-\frac{t^{2}\Sigma^{2}}{2}+O(t^{3})$ where
@@
The proof of Theorem 2.25 proceeds similarly to the proof of [39, Theorem] and it is based on Taylor expansion of order 22 of the functions $\ln({\lambda}_{j}(z))$ around the origin. The exact details are left for the reader.

### v2

@@ 3.1.2. A truncation argument
Let Assumption 2.7 hold with $a<\infty$ and let $b$ be defined by $1/k_{0}=1/a+1/b$ (note that $b=\frac{3k_{0}}{a-k_{0}}$). Let $\zeta,\varepsilon,c_{0},c_{1}$ be like in that assumption. Let us fix some $M>0$. Define a function $G_{M}:{\mathbb{R}}\to{\mathbb{R}}$ as follows. Set $G_{M}(x)=x$ if $|x|\leq M$, set $G_{m}(x)=0$ if $|x|\geq 2M$ and on [2M,M][-2M,-M],  let $G_{M}$ identify with the linear function connecting (2M,0)(-2M,0) and (M,M)(-M,-M), while on $[M,2M]$ let it identify with the linear function connecting $(2M,0)$ and $(M,M)$. Then
@@ 3.2.2. A truncation argument
where $1/k_{0}=1/p+1/a$ for $a\leq p$. Therefore, with $\tilde{S}_{n}=\sum_{j=1}^{k_{n}}\tilde{Y}_{j,n}$ we have
(3.10) $\|S_{n}-\tilde{S}_{n}\|_{L^{k_{0}}}\leq C(1+\sigma_{n}^{2}M_{n}^{-p/a})=O(1)$
@@ 3.4. Reduction of optimal CLT rates from two sided functionals to one sided ones
Finally to deduce Theorem 2.22 for $S_{n}f$ from the corresponding result for $S_{n}g$ let us take some b<a1b<a-1. By Theorem 2.22, we can couple $S_{n}g$ with a standard normal random variable $Z$ so that $\|S_{n}g/{\sigma}_{n}-Z\|_{L^{b}}\leq C{\sigma}_{n}^{-(1-u)}$. Now by Berkes–Philipp Lemma [6, Lemma A.1], we can also couple all three random variables $S_{n}g$, $S_{n}f$ and $Z$ so that (3.17) still holds under the new probability law. ∎
@@ 4.1. Products of positive matrices and other operators
We assume that there are (possibly random) Birkhoff cones ${\mathcal{C}}_{j}\subset B_{j}$ and $n_{0}\in{\mathbb{N}}$ such that for all $j$ we have $A(X_{j}){\mathcal{C}}_{j}\subset{\mathcal{C}}_{j+1}$ (almost surely) and the projective diameter of $\textbf{A}_{j}^{n_{0}}{\mathcal{C}}_{j}$ inside ${\mathcal{C}}_{j+n_{0}}$ does not exceed some constant $d_{0}<\infty$ which is independent of $j$ (see [39, Appendix A] for an overview of projective metrics and cones). Finally, let us assume that the cones are regenerating in the sense of [63, Section 5], namely there exist $r\in{\mathbb{N}}$ and $C>0$ such that every $g\in B_{j}$ can be written as
@@
Each $A_{j}(X_{j})$ is the random transfer operator associated with a random expanding map $T_{X_{j}}$ satisfying the conditions of [39, Ch.5]. That is
@@
for a Hölder continuous random functions $\phi_{X_{j}}$ (which are uniformly Hölder continuous) and Hölder continuous functions $g$. Then by [39, Theorem] there are random cones satisfying the above conditions and (4.1) holds.
@@ 4.4.1. Processes with initial condition
First, by [20, Corollary 5.3] we have
@@
where the second inequality uses the Hölder inequality. Notice next that by applying again [20, Corollary 5.3] we see that
@@
for all $s\geq 0$ and measurable sets $A\in\sigma\{X_{j}:j\leq s\}$ and $B\in{\sigma}\{X_{j}:j>s\}$. Note that in the notations of [10, Eq. (1.6) and Eq. (2.2)] we have $\psi_{U}(1)=\psi^{*}(1)-1$.
@@ 4.4.2. Processes without initial condition
Now, by applying again [20, Corollary 5.3] we see that
@@ 4.5. Application to $\text{GARCH}(\mathfrak{p},\mathfrak{q})$ sequences
with $\mu,\alpha_{i},\beta_{j}\in{\mathbb{R}},\mu>0$. We refer to [36, Example 3.5] for more references and motivation for considering such processes. We assume here that with $r=\max(\mathfrak{p},\mathfrak{q})$,
@@
Now, as explained in [36, Example 3.5] we have
@@
Arguing like in [36, Example 3.5] one can show that
@@ 4.8. Subshifts of finite type and Hölder on average observables
Another example which is relevant to our setup is when working with measure of maximal entropy of a subshift of finite type $T$. Let us briefly recall the definition of a subshift of finite type. Let $\mathcal{A}$} be a finite set and let $(A_{i,j})_{i,j\in{\mathcal{A}}}$ be a matrix with 010-1 entries such that $A^{M}$ has only positive entries for some $M$. Let $\Sigma=\{(x_{i})\in{\mathcal{A}}^{{\mathbb{N}}}:A_{x_{i},x_{i+1}}=1\}$ and let $\sigma:\Sigma\to\Sigma$ be the left shift. Let $\mu$ be the unique measure of maximal entropy (see [7]). Then (see [7]) when viewed as random variables $(X_{j})_{j\geq 0}$ whose path is distributed according to $\mu$ the coordinates $X_{j}$ form a $\psi$-mixing Markov chain, and so we can prove optimal CLT rates for partial sums of the form $\sum_{j=0}^{n-1}f_{j}\circ T^{j}$ for uniformly bounded functions $f_{j}$ which are only Hölder continuous on average, that is under the assumption that
@@
for some $s$ and $\delta$. Indeed, Assumption 2.6 is in force. Note that we can also consider Markov measures on Gibbs-Markov maps [2] since also in that setup the coordinates are $\psi$-mixing exponentially fast. When the functions are not uniformly bounded but instead are bounded in some $L^{p}$ norm then we get rates ${\sigma}_{n}^{1-u}$ where $u\to 0$ as $p\to\infty$. In fact, even when $f_{j}=f$ does not depend on $j$ these results seem to be new. Moreover, we can also consider non-stationary SFT $(T_{j})$ like in [25] since also in that case the coordinates are $\psi$-mixing (see [25, 40]). This allows us to prove optimal CLT rates for non-stationary Markovian piecewise expanding intervals maps and Hölder on average functions $f_{j}$, see [25, Section 4].
Another application is to Gibbs Markov maps considered in [2].
@@
When considering non stationary SFT (see [25]) we can get results for Markov measures, and so also for small perturbations of Anosov maps, see [25, Appendix C].
@@ 5.4. A complex Perron Frobenius theorem
Under one of Assumptions 2.6, 2.7 and 2.10 denote $B_{j}={\mathcal{B}}_{j,q,p,\delta^{\prime}}$ (where under Assumption 2.6 we have $\delta^{\prime}=\delta$). Under Assumption 2.11 denote $B_{j}={\mathcal{B}}_{j,\infty,\infty,\delta^{1/2}}$. Next, by applying [25, Theorem D.2] we get the following corollary of Theorem 5.1.
@@ 6.1. A martingale coboundary representations, the asymptotic behavior of the variance
Finally, the proof of the last statement proceeds like the proof of [4, Theorem 3.5], with minor modifications. ∎
@@ 6.2. Quadratic variation and moment estimates
First, to simplify the notation let us assume that $j=0$. The argument below is similar to the first part of the proof of [16, Theorem 4.1]. First, we write
@@
for some constant $C_{q}>0$. Recall the following version of Burkholder’s inequality for martingales (see [55, Theorem 2.12]). Let $\mathfrak{d}_{1},....,\mathfrak{d}_{n}$ be a martingale difference with respect to a filtration $({\mathcal{G}}_{j})_{j=1}^{n}$ on a probability space. Let $D_{n}=\mathfrak{d}_{1}+\mathfrak{d}_{2}+...+\mathfrak{d}_{n}$ and $E_{n}=\mathfrak{d}_{1}^{2}+\mathfrak{d}_{2}^{2}+...+\mathfrak{d}_{n}^{2}$. Then, for every $s\geq 2$ there are constants $c_{s},C_{s}>0$ depending only on $s$ such that
@@ 6.6. The sequential pressure function and its approximation properties
There exist $r_{0}>0,C_{0}>0$ and $n_{0}$ such that on $t\in[-r_{0},r_{0}]$ for all $n\geq n_{0}$ we can develop a branch $\Lambda_{j,n}(t)$ of $\ln\mu_{j}(e^{itS_{j,n}g})={\mathbb{E}}[e^{itS_{j,n}g}]$ such that for s=0,1,2,3,,ks=0,1,2,3,...,k we have
@@
Notice that the first $k$ derivatives of $t\to\bar{E}_{j}^{t,n}$ are also of order $O(\delta_{2}^{n})$ (by the differentiation rule for derivatives of products). Hence for s=0,1,2,3,,ks=0,1,2,3,...,k we have
@@
for some constant $C^{\prime}_{k}$. Thus Theorems 2.21 and 2.22 follow from [41, Theorem 5] and [41, Theorem 9] and [41, Corollary 11].
@@
As before, by Proposition 3.9 it is enough to prove both theorems for $S_{n}g$ (or $\tilde{S}_{n}$ in the notations of Section 3.2). Let $\Lambda_{j,n}(t)$ be the branch of $\ln{\mathbb{E}}[e^{itS_{j,n}g}]$ from Lemma 6.5. Note also that by rescaling under one of Assumptions 2.7 or 2.10 we can always assume that $u=0$ in Theorems 2.21 and 2.22. Recall that we assumed that ${\mathbb{E}}[f_{j}]=0$ which implies that ${\mathbb{E}}[g_{j}]=0$. By applying again [41, Theorem 5] and [41, Theorem 9] and [41, Corollary 11]. Theorems 2.21 and 2.22 for $S_{n}g$ will follow from the following result.
@@
In the circumstances of Theorems 2.21 and 2.22 there are constants ${\delta}_{k}>0$ and $C_{k}>0$ such that for all $s\leq k$ we have
(6.11) $\sup_{t\in[-{\delta}_{k},{\delta}_{k}]}|\Lambda^{(s)}_{0,n}(t)|\leq C_{3}{\sigma}_{n}^{2}.$
@@
Notice that it is enough to prove (6.11) with $s=k$. Indeed, if $|t|\leq\delta_{3}$ then
$\Lambda^{(k-1)}_{0,n}(t)=\int_{0}^{t}\Lambda^{(k)}_{0,n}(x)dx=O({\sigma}_{n}^{2})$
and similarly $\Lambda^{(s)}_{0,n}(t)=O({\sigma}_{n}^{2})$ for s<k1s<k-1.
@@
Applying [24, Lemma 43] with $S=S_{j,n}g=S_{j,n}g-{\mathbb{E}}[S_{j,n}g]$ we see that there is $r=r(B)$ such that if $t\in[-r,r]$ then
(6.13) $|\Lambda_{j,n}^{(k_{0})}(t)|\leq D_{0}{\mathbb{E}}[|S|^{k_{0}}]$
for some constant $D_{0}$. Now the result follows since $\|S_{j,n}f-S_{j,n}g\|_{L^{k_{0}}}\leq C$ for some constant $C$ (see (3.10), (3.6) and Lemma 3.5). ∎
@@
$\sup_{t\in[-r_{0},r_{0}]}\left|\Pi_{I_{j}}^{(k_{0})}(t)\right|\leq D_{0}(1+{\mathbb{E}}[|S_{I_{j}}f|^{k_{0}}]).$
@@
(6.16) $\sup_{t\in[-r_{k},r_{k}]}\left|\Pi_{0,n}^{(k_{0})}(t)\right|\leq D_{0}k_{n}+\sum_{j=1}^{k_{n}}{\mathbb{E}}[|S_{I_{j}}f|^{k_{0}}]=O({\sigma}_{n}^{2})$
@@
$\sup_{t\in[-r_{0},r_{0}]}\left|\Lambda_{n}^{(k_{0})}(t)\right|\leq\tilde{D}{\sigma}_{n}^{2}$
@@ 7. Large deviations
where $g_{j}$ are the functions from Lemma 3.5. Denote $B_{j}={\mathcal{B}}_{j,\infty,\infty,\delta^{1/2}}$. Then, since $\sup_{j}\|g_{j}\|_{j,\infty,\infty,\delta^{1/2}}<\infty$ we see that $L_{j,z}$ are uniformly analytic in $z$. By applying [25, Theorem D.2] we get the following corollary of Theorem 5.1.
@@
Now, notice that (see the arguments in [39, Ch.5]) $\int\ln{\lambda}_{\omega}(t)d\mathbb{P}(\omega)=1-\frac{t^{2}\Sigma^{2}}{2}+O(t^{3})$ where
@@
The proof of Theorem 2.25 proceeds similarly to the proof of [40, Theorem] and it is based on Taylor expansion of order 22 of the functions $\ln({\lambda}_{j}(z))$ around the origin. The exact details are left for the reader.